# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v1

@@ 3.2 Data Curation and Final Dataset Statistics
We perform manual curation to remove videos with strong artifacts to ensure the quality of the dataset. Common artifact types included unnatural head movements, lip-sync errors, and background warping. Detailed per-generator artifact observations, which inform our curation and highlight unique generator characteristics, are provided in the Supplementary. We then separate the generated data into a train and test split, ensuring identity separation, i.e., the source image and the driving signal used for generating any training deepfake video is not used for generating testing deepfake videos. Tab. 4 details the final statistics and splits, and sample frames are shown in Fig. 1.
Table 4: Summary statistics of the TalkingHeadBench across generators and associated visual artifacts.
Data Type # Total # Train # Test Ratio Removed Key Artifacts
Hallo xu2024hallo 420 303 117 $\sim 42\%$ Lip sync error, hair artifacts, occlusion error, background warping
Hallo2 cui2024hallo2 432 327 105 $\sim 60\%$ Lip sync error, hair artifacts, occlusion error, background warping
AniPortrait (Audio) wei2024aniportrait 542 396 146 $\sim 66.7\%$ Lip sync error, inconsistent head movement, teeth/pupil anomalies
AniPortrait (Video) wei2024aniportrait 422 314 108 $\sim 80\%$ Extreme face/background warping, occlusion errors, hair artifacts
LivePortrait  guo2024liveportrait 529 352 177 $\sim 60\%$ Face distortion, inconsistent head movement, occlusion error
EMOPortraits drobyshev2024emoportraits 573 381 192 $\sim 50\%$ Floating head, edge anomalies, spatial inconsistencies
MAGI-1 magi1 66 0 66 $\sim 18\%$ Face/background warping
TOTAL Fakes 2984 2073 911 $\sim 63\%$
@@ 4.1 Evaluation Setup
(i) How does the performance of the SOTA detectors change from FaceForensics++ rossler2019faceforensics++ , where detectors have near-perfect accuracy, to TalkingHeadBench?
@@
SOTA Detectors To address these questions, we benchmark four publicly available SOTA deepfake detectors: CADDM dong2023implicit , LipFD liu2024lips , DeepFake-Adapter shao2025deepfake , and TALL xu2023tall . Similar to the generators, we aim to diversify the modalities and frameworks of our detectors, as shown in Tab. 3. These four detectors meet our goal of being the latest SOTA detectors with available code, and together, they capture varying modalities.
@@
Evaluation Protocols To systematically evaluate the generalization of these SOTA detectors, we design three distinct protocols:
@@
P1 (Identity Shift Only) Evaluates generalization to unseen identities within familiar generators. Models are trained on deepfakes from all generators and real videos, then tested on each generator and real videos, ensuring no overlap in identities between the training and test sets. (see Tab. 5).
@@
P2 (Generator Shift) Evaluates generalization to an unseen generator. Models are trained on remaining generators and real videos, then tested on the held-out generator and real videos (see Tab. 6).
@@
P3 (Identity & Generator Shift) Evaluates generalization to an unseen generator and identities. Using a leave-one-out approach, models are trained on the train splits of remaining generators and real videos, then tested on the test split of the held-out generator and real videos (see Tab. 7).
@@
For these protocols, we select 1600 real videos from FaceForensics++ rossler2019faceforensics++ and CelebV-HQ zhu2022celebv , unseen during generation, to match fake video numbers per protocol, and perform face comparison to prevent identity leakage between train and test across real and fake video sets.
@@
Evaluation Metrics Detector performance under these protocols is measured with three metrics: AUC, TPR@FPR=1% (T1), and TPR@FPR=0.1% (T0.1), following IJB-C face benchmark maze2018iarpa . Stricter thresholds on FPR, e.g., T0.1, are more useful when these detectors are applied at scale and only a small amount of bad detections are tolerable. We will prioritize T1 for most analysis in this paper, following its use in the IJB-C face benchmark, but later we will analyze the detector performance across various thresholds in FPR. We classify the performance of detectors as:
Excellent: $\geq$0.99, Good: 0.95-0.99, Reasonable: 0.85-0.95, Fair: 0.75-0.85, and Poor: <0.75.
Table 5: Protocol 1 — Identity Shift Only. Best numbers per (generator) are bold.
AniAudio wei2024aniportrait AniVideo wei2024aniportrait Hallo xu2024hallo Hallo2 cui2024hallo2 EMOPortraits drobyshev2024emoportraits LivePortrait guo2024liveportrait
Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1
CADDM dong2023implicit 0.90 0.55 0.25 0.83 0.39 0.39 0.94 0.65 0.64 0.96 0.45 0.45 0.94 0.27 0.23 0.95 0.59 0.42
TALL xu2023tall 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.99 1.00 0.99 0.98 1.00 0.98 0.98 1.00 0.97 0.96
Lip FD liu2024lips 0.98 0.75 0.58 0.99 0.74 0.38 0.98 0.72 0.54 0.99 0.83 0.62 0.99 0.78 0.50 0.98 0.84 0.56
DF-Adapter shao2025deepfake 0.99 0.58 0.06 1.00 1.00 0.15 1.00 0.99 0.11 1.00 0.99 0.98 1.00 0.99 0.25 1.00 0.98 0.41
Table 6: Protocol 2 — Generator Shift Only. Best numbers per (generator) are bold
AniAudio wei2024aniportrait AniVideo wei2024aniportrait Hallo xu2024hallo Hallo2 cui2024hallo2 EMOPortraits drobyshev2024emoportraits LivePortrait guo2024liveportrait
Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1
CADDM dong2023implicit 0.88 0.32 0.02 0.94 0.27 0.18 0.97 0.61 0.52 0.94 0.37 0.14 0.79 0.02 0.00 0.98 0.71 0.63
TALL xu2023tall 1.00 1.00 0.98 1.00 0.99 0.98 1.00 1.00 1.00 0.97 0.89 0.81 0.97 0.70 0.40 1.00 1.00 0.98
LipFD liu2024lips 0.99 0.76 0.51 0.97 0.62 0.30 0.98 0.70 0.47 0.87 0.50 0.32 0.89 0.30 0.08 0.99 0.81 0.63
DF-Adapter shao2025deepfake 1.00 1.00 0.83 1.00 1.00 1.00 1.00 1.00 0.73 0.97 0.80 0.68 0.99 0.93 0.80 1.00 1.00 0.99
Table 7: Protocol 3 — Identity & Generator Shift Combined. Best numbers per (generator) are bold
AniAudio wei2024aniportrait AniVideo wei2024aniportrait Hallo xu2024hallo Hallo2 cui2024hallo2 EMOPortraits drobyshev2024emoportraits LivePortrait guo2024liveportrait
Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1
CADDM dong2023implicit 0.82 0.25 0.21 0.98 0.60 0.60 0.95 0.55 0.20 0.93 0.17 0.17 0.75 0.03 0.03 0.95 0.52 0.41
TALL xu2023tall 1.00 1.00 0.99 1.00 1.00 1.00 1.00 0.99 0.99 0.97 0.84 0.82 0.97 0.48 0.48 1.00 0.93 0.92
LipFD liu2024lips 0.99 0.72 0.62 0.98 0.71 0.56 0.98 0.72 0.34 0.89 0.57 0.40 0.91 0.30 0.14 0.98 0.77 0.61
DF-Adapter shao2025deepfake 1.00 0.93 0.16 1.00 0.99 0.94 1.00 1.00 0.95 0.98 0.88 0.80 0.98 0.70 0.53 1.00 0.95 0.56
4.2 Impact of Generator and Identity Shift on Performance
@@
Protocol-Dependent Performance Shifts: Overall, detectors exhibit performance shifts across protocols. In particular, average $AUC$ and T1 scores tend to decline from P1 to P2, indicating that generator shift (P2) poses greater challenges than identity shift (P1). This trend continues in P3, where combined generator and identity shifts further reduce performance as illustrated in Fig. 2a. For instance, TALL’s xu2023tall average T1 score decreases from 0.99 in P1 to 0.93 in P2 and 0.87 in P3. Despite these increasingly difficult conditions, TALL and DeepFake-Adapter shao2025deepfake maintain strong performance across all three protocols, with DeepFake-Adapter’s lower score on P1 largely due to poor performance on AniPortrait Audio wei2024aniportrait (T1 = 0.58; see Tab. 5). This underscores the need to assess detectors not just by averages, but also by generator-level sensitivity.
Generator Difficulty and Cross-Protocol Generalization: Fig. 2c reveals the variation of detector performance across generators. EMOPortraits drobyshev2024emoportraits and Hallo2 cui2024hallo2 emerge as the most challenging generators, with average T1 scores below 0.6 and 0.7, respectively. In contrast, LivePortrait guo2024liveportrait and Hallo xu2024hallo consistently yield higher detection rates, with average T1 over 0.8. Notably, while EMOPortraits drobyshev2024emoportraits achieves a high average AUC of $\geq$ 0.90 across all 3 protocols, its T1 drops sharply from 0.80 in P1 to 0.49 in P2, and further to 0.38 in P3. This disparity underscores a critical limitation in current detectors: although models may appear performant under AUC, they struggle to maintain true positive rates under stricter false positive constraints, especially when facing generator and identity shifts.
@@
Figure 2: Average performance scores across detectors and generators. (a) shows the average T1 across all detectors for each evaluation protocol. Detector performance consistently declines from P1 to P2, and further in P3, highlighting the increasing difficulty of generalization. These results suggest that evaluating detectors solely on average performance may obscure generator-specific weaknesses. (b) presents the average TPR at varying FPR thresholds for TALL and DeepFake-Adapter. TALL maintains high TPR even at extreme thresholds (e.g., FPR=0.001), demonstrating strong robustness. In contrast, DeepFake-Adapter exhibits a sharp performance drop as the threshold tightens, highlighting its reduced reliability under stricter operating conditions. (c) shows the average T1 score per generator across all detectors. EMOPortraits and Hallo2 emerge as the most challenging generators, while Hallo and LivePortrait are comparatively easier to detect.
@@ 4.3 Detector Robustness and SOTA Evaluation
Overall Robustness Profile: While detailed per protocol analysis is valuable, evaluating robustness across thresholds provides a comprehensive view of a detector’s generalization capabilities. Fig. 3 summarizes performance across all generators, averaged over protocols, for T1 and T0.1 thresholds. High and balanced coverage across all axes indicates consistent generalization across identity, generator, and joint shifts—crucial for handling unseen data from future SOTA generators and varying identities. At T1 (Fig. 3a), DeepFake-Adapter shao2025deepfake and TALL xu2023tall exhibit strong performance across most generators, with DeepFake-Adapter excelling on LivePortrait guo2024liveportrait , Hallo xu2024hallo , and AniPortraitVideo wei2024aniportrait . However, when the threshold is tightened to T0.1 (Fig. 3b), DeepFake-Adapter’s performance drops sharply—particularly for AniPortraitAudio wei2024aniportrait and EMOPortraits drobyshev2024emoportraits —indicating reduced robustness under stricter constraints. In contrast, TALL retains high scores across most generators at T0.1, suggesting better threshold-level resilience and stronger generalization under a low false positive rate.
Performance across different FPR thresholds: While our primary evaluation focuses on T1 due to its importance in deepfake detection scenarios, analyzing stricter thresholds such as T0.1, T0.01, T0.001 also provides valuable insights into overall detector reliability. As shown in Fig. 2b, even though both TALL and DeepFake-Adapter achieve $\geq$ 0.90 for T1, while FPR decreases, we notice a significant drop for DeepFake-Adapter. This indicates its reduced reliability in real-world applications where minimizing such false accusations is paramount. Meanwhile, TALL demonstrates robust performance, maintaining a consistently high true positive rate even at these very low FPRs, highlighting its suitability for deployment in sensitive environments.
Figure 3: Detector performance across generators measured by T1 and T0.1, averaged over all protocols. (a) shows T1 performance, where DeepFake-Adapter and TALL demonstrate strong generalization across identity, generator, and joint shifts. (b) shows results at a stricter threshold (T0.1), where most detectors have a noticeable drop, highlighting challenges in maintaining high recall under low FPR. Despite this, TALL remains the most robust, showing consistent performance across generators even under tighter operating conditions.

### v2

@@ 3.2 Data Curation and Final Dataset Statistics
In order to create a high-quality dataset, we used a five-stage curation pipeline conducted by experts on our team. This process removed approximately 63% of the nearly 8,000 videos we initially generated, ensuring the final dataset is not easily solved by detectors that rely on simple artifacts.
  1. 1.
    Generator-Specific Artifact Definition: We analyzed videos from each generator to create a taxonomy of common artifacts, such as severe background warping or unnatural static hair. This formed a consistent set of rejection criteria, which are detailed in the supplementary.
  2. 2.
    Primary Review Pass: A curator reviewed all videos from a specific generator, removing samples that matched the artifact definitions or had other global distortions.
  3. 3.
    Cross-Validation by Experts: To ensure consistency, two other curators independently reviewed the remaining videos and removed any additional low-quality samples. Disagreements were resolved through discussion.
  4. 4.
    Detector-Based Final Checkpoint: To further refine the dataset’s difficulty, we passed the remaining videos through four pre-trained SOTA detectors. Any video detected with high confidence ($\geq$ 90% fake probability) across all four detectors was removed. This step filtered out samples that, while appearing plausible to humans, contained subtle but easily machine-detectable cues.
  5. 5.
    Identity-Separated Splits: The final curated set was partitioned into training and testing splits, with strict separation enforced by disallowing any overlap in source image-driving signal pairings across splits. The final dataset statistics are detailed in Tab. 4.
@@ 4.1 Evaluation Setup
Table 5: Overall benchmark. Three protocols reported as stacked subtables. Best numbers per generator in bold.
(a) Protocol 1 — Identity Shift Only.
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.90 0.55 0.25 0.17 0.83 0.39 0.39 0.21 0.94 0.65 0.64 0.16 0.96 0.45 0.45 0.11 0.94 0.27 0.23 0.17 0.95 0.59 0.42 0.13 HiFi-Net [hifi_net_xiaoguo] 0.95 0.47 0.27 0.09 0.95 0.66 0.41 0.09 0.93 0.39 0.19 0.11 0.92 0.22 0.12 0.12 0.81 0.03 0.01 0.25 0.90 0.26 0.04 0.12 AltFreezing [wang2023altfreezing] 1.00 0.93 0.71 0.03 1.00 0.91 0.75 0.05 1.00 0.97 0.97 0.03 1.00 0.95 0.88 0.05 1.00 0.93 0.89 0.04 1.00 0.94 0.89 0.04 TALL [xu2023tall] 1.00 1.00 1.00 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 1.00 0.99 0.98 0.01 1.00 0.98 0.98 0.05 1.00 0.97 0.96 0.01 LipFD [liu2024lips] 0.98 0.75 0.58 0.05 0.99 0.74 0.38 0.05 0.98 0.72 0.54 0.05 0.99 0.83 0.62 0.04 0.99 0.78 0.50 0.05 0.98 0.84 0.56 0.05 DF-Adapter [shao2025deepfake] 0.99 0.58 0.06 0.01 1.00 1.00 0.15 0.01 1.00 0.99 0.11 0.01 1.00 0.99 0.98 0.01 1.00 0.99 0.25 0.01 1.00 0.98 0.41 0.01 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.97 0.65 0.26 0.14 0.97 0.55 0.19 0.15 0.97 0.63 0.26 0.14 0.96 0.57 0.21 0.15 0.93 0.34 0.11 0.14 0.91 0.28 0.04 0.15
(b) Protocol 2 — Generator Shift Only.
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.88 0.32 0.02 0.18 0.94 0.27 0.18 0.09 0.97 0.61 0.52 0.08 0.94 0.37 0.14 0.12 0.79 0.02 0.00 0.21 0.98 0.71 0.63 0.06 HiFi-Net [hifi_net_xiaoguo] 0.88 0.08 0.01 0.16 0.82 0.16 0.06 0.18 0.82 0.02 0.00 0.20 0.48 0.02 0.00 0.45 0.58 0.01 0.00 0.34 0.85 0.01 0.01 0.20 AltFreezing [wang2023altfreezing] 1.00 0.97 0.82 0.02 1.00 0.94 0.66 0.02 1.00 0.98 0.87 0.01 0.98 0.71 0.53 0.06 0.96 0.69 0.52 0.08 1.00 0.94 0.78 0.02 TALL [xu2023tall] 1.00 1.00 0.98 0.01 1.00 0.99 0.98 0.02 1.00 1.00 1.00 0.00 0.97 0.89 0.81 0.08 0.97 0.70 0.40 0.15 1.00 1.00 0.98 0.01 LipFD [liu2024lips] 0.99 0.76 0.51 0.05 0.97 0.62 0.30 0.09 0.98 0.70 0.47 0.08 0.87 0.50 0.32 0.17 0.89 0.30 0.08 0.23 0.99 0.81 0.63 0.05 DF-Adapter [shao2025deepfake] 1.00 1.00 0.83 0.01 1.00 1.00 1.00 0.00 1.00 1.00 0.73 0.01 0.97 0.80 0.68 0.15 0.99 0.93 0.80 0.13 1.00 1.00 0.99 0.01 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.93 0.39 0.04 0.12 0.95 0.36 0.14 0.11 0.94 0.34 0.08 0.15 0.97 0.51 0.25 0.06 0.89 0.15 0.07 0.06 0.88 0.23 0.11 0.16
(c) Protocol 3 — Identity & Generator Shift Combined.
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.82 0.25 0.21 0.023 0.98 0.60 0.60 0.06 0.95 0.55 0.20 0.11 0.93 0.17 0.17 0.10 0.75 0.03 0.03 0.24 0.95 0.52 0.41 0.09 HiFi-Net [hifi_net_xiaoguo] 0.88 0.03 0.01 0.17 0.86 0.27 0.11 0.16 0.84 0.00 0.00 0.18 0.48 0.00 0.00 0.44 0.61 0.01 0.00 0.34 0.84 0.01 0.00 0.27 AltFreezing [wang2023altfreezing] 1.00 0.87 0.78 0.03 1.00 0.94 0.88 0.03 1.00 0.95 0.89 0.02 0.99 0.85 0.64 0.04 0.95 0.58 0.26 0.08 0.99 0.90 0.56 0.03 TALL [xu2023tall] 1.00 1.00 0.99 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 0.97 0.84 0.82 0.06 0.97 0.48 0.48 0.15 1.00 0.93 0.92 0.02 LipFD [liu2024lips] 0.99 0.72 0.62 0.04 0.98 0.71 0.56 0.08 0.98 0.72 0.34 0.08 0.89 0.57 0.40 0.15 0.91 0.30 0.14 0.22 0.98 0.77 0.61 0.06 DF-Adapter [shao2025deepfake] 1.00 0.93 0.16 0.02 1.00 0.99 0.94 0.01 1.00 1.00 0.95 0.00 0.98 0.88 0.80 0.09 0.98 0.70 0.53 0.14 1.00 0.95 0.56 0.03 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.94 0.49 0.06 0.13 0.96 0.40 0.16 0.12 0.93 0.33 0.06 0.18 0.98 0.55 0.27 0.06 0.87 0.06 0.03 0.18 0.85 0.08 0.03 0.17
@@
(i) How does the performance of the SOTA detectors change from FaceForensics++ [rossler2019faceforensics++], where detectors have near-perfect accuracy, to TalkingHeadBench?
@@
SOTA Detectors To address these questions, we benchmark seven publicly available SOTA deepfake detectors: CADDM [dong2023implicit], LipFD [liu2024lips], DeepFake-Adapter [shao2025deepfake], TALL [xu2023tall], AltFreezing [wang2023altfreezing], MM-Det [song2025learningmultimodalforgeryrepresentation], and HiFi-Net [hifi_net_xiaoguo]. Similar to the generators, we aim to diversify the modalities and frameworks of our detectors, as shown in Tab. 3. These seven detectors meet our goal of being the latest SOTA detectors with available code, and together, they capture varying modalities.
@@
Evaluation Protocols To systematically evaluate the generalization of these SOTA detectors, we design three protocols:
@@
P1 (Identity Shift Only) Evaluates generalization to unseen identities within familiar generators. Models are trained on deepfakes from all generators and real videos, then tested on each generator and real videos, ensuring no overlap in identities between the training and test sets. (see Tab. 5(a)).
@@
P2 (Generator Shift) Evaluates generalization to an unseen generator. Models are trained on remaining generators and real videos, then tested on the held-out generator and real videos (see Tab. 5(b)).
@@
P3 (Identity & Generator Shift) Evaluates generalization to an unseen generator and identities. Using a leave-one-out approach, models are trained on the train splits of remaining generators and real videos, then tested on the test split of the held-out generator and real videos (see Tab. 5(c)).
@@
For these protocols, we select 1600 real videos from FaceForensics++ [rossler2019faceforensics++] and CelebV-HQ [zhu2022celebv], unseen during generation, to match fake video numbers per protocol, and perform face comparison to prevent identity leakage between train and test across real and fake video sets.
@@
Evaluation Metrics Detector performance under these protocols is measured with three metrics: AUC, Brier Score, TPR@FPR=1% (T1), and TPR@FPR=0.1% (T0.1), following IJB-C face benchmark [maze2018iarpa]. Stricter thresholds on FPR, e.g., T0.1, are more useful when these detectors are applied at scale and only a small amount of bad detections are tolerable. We will prioritize T1 for most analysis in this paper, following its use in the IJB-C face benchmark, but later we will analyze the detector performance across various thresholds in FPR. We classify the performance of detectors as:
Excellent: $\geq$0.99, Good: 0.95-0.99, Reasonable: 0.85-0.95, Fair: 0.75-0.85, and Poor: <<0.75.
4.2 Impact of Generator and Identity Shifts
@@
Protocol-Dependent Performance Shifts: Overall, detectors exhibit performance shifts across protocols. In particular, average $AUC$ and T1 scores tend to decline from P1 to P2, indicating that generator shift (P2) poses greater challenges than identity shift (P1). This trend continues in P3, where combined generator and identity shifts further reduce performance as illustrated in Fig. 2a. For instance, TALL’s [xu2023tall] average T1 score decreases from 0.99 in P1 to 0.93 in P2 and 0.87 in P3. Despite these increasingly difficult conditions, TALL and DeepFake-Adapter [shao2025deepfake] maintain strong performance across all three protocols, with DeepFake-Adapter’s lower score on P1 largely due to poor performance on AniPortrait Audio [wei2024aniportrait](T1 = 0.58; see Tab. 5(a)). This underscores the need to assess detectors not just by averages, but also by generator-level sensitivity.
Generator Difficulty and Cross-Protocol Generalization: Fig. 2b shows the variation of detector performance across generators. EMOPortraits [drobyshev2024emoportraits] stands out as the most challenging case, with average T1 of only 0.44, far below all others. By contrast, the remaining generators cluster within a narrower range of 0.60-0.69, indicating more comparable levels of detectability. This sharp gap emphasizes that while detectors maintain moderate robustness across most generators, EMOPortraits remains a persistent failure case that reveals their limitations under stricter thresholds.
@@
Figure 2: Average performance scores across detectors and generators. (a) shows the average T1 across all detectors for each evaluation protocol. Detector performance consistently declines from P1 to P2, and further in P3, highlighting the increasing difficulty of generalization. These results suggest that evaluating detectors solely on average performance may obscure generator-specific weaknesses. (b) shows the average T1 score per generator across all detectors. EMOPortraits emerges as the most challenging generator, while the others fall within a narrower range of 0.60–0.69, indicating relatively comparable levels of detectability. (c) presents the average TPR at varying FPR thresholds for TALL and DeepFake-Adapter. TALL maintains high TPR even at extreme thresholds (e.g., FPR=0.001), demonstrating strong robustness. In contrast, DeepFake-Adapter exhibits a sharp performance drop as the threshold tightens, highlighting its reduced reliability under stricter operating conditions.
@@ 4.3 Detector Robustness and SOTA Evaluation
Overall Robustness Profile: While detailed per protocol analysis is valuable, evaluating robustness across thresholds provides a comprehensive view of a detector’s generalization capabilities. Fig. 3 summarizes performance across all generators, averaged over protocols, for T1 and T0.1 thresholds. High and balanced coverage across all axes indicates consistent generalization across identity, generator, and joint shifts—crucial for handling unseen data from future SOTA generators and varying identities. At T1 (Fig. 3a), DeepFake-Adapter [shao2025deepfake] and TALL [xu2023tall] exhibit strong performance across most generators, with DeepFake-Adapter excelling on LivePortrait [guo2024liveportrait], Hallo [xu2024hallo], and AniPortraitVideo [wei2024aniportrait]. However, when the threshold is tightened to T0.1 (Fig. 3b), DeepFake-Adapter’s performance drops sharply—particularly for AniPortraitAudio [wei2024aniportrait] and EMOPortraits [drobyshev2024emoportraits]—indicating reduced robustness under stricter constraints. In contrast, TALL retains high scores across most generators at T0.1, suggesting better threshold-level resilience and stronger generalization under a low false positive rate. We also evaluate detector performance generalizing to unseen generators such as commercial generator MAGI-1 [magi1] and academic generator Hallo3 [cui2024hallo3]. We observe similar trend whereas the best-performing models, TALL and DeepFake-Adapter, maintain high detection scores. We include the full result in the Supplementary.
Performance across different FPR thresholds: While our primary evaluation focuses on T1 due to its importance in deepfake detection scenarios, analyzing stricter thresholds such as T0.1, T0.01, T0.001 also provides valuable insights into overall detector reliability. As shown in Fig. 2c, even though both TALL and DeepFake-Adapter achieve $\geq$ 0.90 for T1, while FPR decreases, we notice a significant drop for DeepFake-Adapter. This indicates its reduced reliability in real-world applications where minimizing such false accusations is paramount. Meanwhile, TALL demonstrates robust performance, maintaining a consistently high true positive rate even at these very low FPRs, highlighting its suitability for deployment in sensitive environments.
Figure 3: Detector performance across generators measured by T1 and T0.1, averaged over all protocols. (a) shows T1 result, where DeepFake-Adapter and TALL demonstrate strong generalization across identity, generator, and joint shifts. (b) shows stricter T0.1 result, where most detectors have a noticeable drop, highlighting challenges in maintaining high recall under low FPR. Despite this, TALL remains the most robust, showing consistent performance across generators even under tighter operating conditions.
4.4 Generalization to Emerging Generators
To simulate a real-world scenario, we evaluated detectors against deepfakes from emerging academic generator Hallo3 [cui2024hallo3] and commercial generator MAGI-1 [magi1], which were not seen during training. The results, provided in supplementary (Section 4), show that top-performing models like TALL and AltFreezing maintained high accuracy, suggesting that training on the diverse and curated set in TalkingHeadBench enables detectors to handle novel generator architectures. This indicates the detectors have learned a fundamental representation of talking-head artifacts beyond overfitting to specific generator fingerprints.